# SpiLiFormer: Enhancing Spiking Transformers with Lateral Inhibition

### v1

@@ Abstract
Spiking Neural Networks (SNNs) based on Transformers have garnered significant attention due to their superior performance and high energy efficiency. However, the spiking attention modules of most existing Transformer-based SNNs are adapted from those of analog Transformers, failing to fully address the issue of over-allocating attention to irrelevant contexts. To fix this fundamental yet overlooked issue, we propose a Lateral Inhibition-inspired Spiking Transformer (SpiLiFormer). It emulates the brain’s lateral inhibition mechanism, guiding the model to enhance attention to relevant tokens while suppressing attention to irrelevant ones. Our model achieves state-of-the-art (SOTA) performance across multiple datasets, including CIFAR-10 (+0.45%), CIFAR-100 (+0.48%), CIFAR10-DVS (+2.70%), N-Caltech101 (+1.94%), and ImageNet-1K (+1.6%)  [^1] These results are obtained by comparing with state-of-the-art spiking neural network models that have a similar number of parameters.. Notably, on the ImageNet-1K dataset, SpiLiFormer (69.9M parameters, 4 time steps, 384 resolution) outperforms E-SpikeFormer (173.0M parameters, 8 time steps, 384 resolution), a SOTA spiking Transformer, by 0.46% using only 39% of the parameters and half the time steps. Our code and training checkpoints will be released upon acceptance.
@@ 1 Introduction
Figure 1: Illustration of the severe attention distraction phenomenon in Transformer-based SNNs, the architecture of the Lateral Inhibition-inspired Spiking Transformer (SpiLiFormer), and the detailed processing mechanism of Feedfoward-pathway Laternal Differential Inhibition (FF-LiDiff) attention and Feedback-pathway Lateral Difference Inhibition (FB-LiDiff) attention. (a) refers to the attention distraction in mainstream Transformer-based SNNs, where models excessively focus on irrelevant background information, leading to misclassification. (b) denotes the retinal lateral inhibition mechanism, illustrating how horizontal cells regulate neural responses to enhance contrast and suppress noise. (c) presents the architecture of SpiLiFormer. (d) and (e) represent the information processing flows of FF-LiDiff and FB-LiDiff attention, respectively.
@@ 3.4 Feedback-pathway Lateral Differential Inhibition Attention
The second forward propagation begins by taking the processed forward representation $S_{\text{stage3}}^{Ds}$ and the corresponding feedback information $X_{\text{FB}}^{1}$ as input to the first FB-LiDiff attention. The objective is to reduce model energy consumption by eliminating redundant computations in the non-feedback loop (i.e., Stage 1 and Stage 2). A detailed energy consumption analysis can be found in  Sec. A.3. Additionally, during the second forward propagation, we share all parameter weights across the FB-LiDiff blocks to minimize the number of model parameters. Given the input $X_{2}^{m}$, the process described above can be formulated as:
@@ 3.5 Training Strategy and Loss Function
where $\alpha$ is a hyperparameter that dynamically balances the losses from two forward propagations. Unless explicitly specified otherwise, $\alpha$ is set to 0.5 in the subsequent experiments. Additionally, $y$ represents the true label; $O_{\text{stage3}}^{N_{3}}$ and $(O_{\text{stage3}}^{N_{3}})^{{}^{\prime}}$ denote the outputs from the first and second forward propagations of the last FB-LiDiff block, respectively; $CH(\cdot)$ represents the classification head module, as described in previous studies [36, 35]; and $\mathcal{L}_{CE}$ denotes the cross-entropy loss function.
@@ 4 Experiments
Dataset Methods Architecture Param(M) Time Step Top-1 Acc(%)
@@ 4.5 Ablation Study
We conduct an ablation study on CIFAR-10 (static) and CIFAR10-DVS (neuromorphic) to assess the impact of FF-LiDiff and FB-LiDiff attention on model performance. All ablation experiments follow the training details in Sec. 4.2 and Sec. 4.3, unless otherwise specified.
@@
As shown in Tab. 3, we observe that the model exhibits varying degrees of performance degradation on both datasets under the two aforementioned scenarios. Specifically, performance exhibits a decrease of 0.31% and 0.62% on the CIFAR-10 dataset, and 1.6% and 3.0% on the CIFAR10-DVS dataset. These results highlight the greater importance of FB-LiDiff attention in the network structure. Furthermore, our analysis indicates that the combination of FF-LiDiff and FB-LiDiff attention helps mitigate the distraction problem more effectively, thereby further improving the model’s performance.
Dataset Methods Top-1 Acc(%)
@@ A.3 Energy Consumption Calculation of SNNs and ANNs
A.4 Supplementary Tables and Figures

### v2

@@ Abstract
Spiking Neural Networks (SNNs) based on Transformers have garnered significant attention due to their superior performance and high energy efficiency. However, the spiking attention modules of most existing Transformer-based SNNs are adapted from those of analog Transformers, failing to fully address the issue of over-allocating attention to irrelevant contexts. To fix this fundamental yet overlooked issue, we propose a Lateral Inhibition-inspired Spiking Transformer (SpiLiFormer). It emulates the brain’s lateral inhibition mechanism, guiding the model to enhance attention to relevant tokens while suppressing attention to irrelevant ones. Our model achieves state-of-the-art (SOTA) performance across multiple datasets, including CIFAR-10 (+0.45%), CIFAR-100 (+0.48%), CIFAR10-DVS (+2.70%), N-Caltech101 (+1.94%), and ImageNet-1K (+1.6%)  [^1] These results are obtained by comparing with state-of-the-art spiking neural network models that have a similar number of parameters.. Notably, on the ImageNet-1K dataset, SpiLiFormer (69.9M parameters, 4 time steps, 384 resolution) outperforms E-SpikeFormer (173.0M parameters, 8 time steps, 384 resolution), a SOTA spiking Transformer, by 0.46% using only 39% of the parameters and half the time steps. The code and model checkpoints are publicly available at https://github.com/KirinZheng/SpiLiFormer.
@@ 1 Introduction
Figure 1: Illustration of the severe attention distraction phenomenon in Transformer-based SNNs, the architecture of the Lateral Inhibition-inspired Spiking Transformer (SpiLiFormer), and the detailed processing mechanism of Feedfoward-pathway Lateral Differential Inhibition (FF-LiDiff) attention and Feedback-pathway Lateral Difference Inhibition (FB-LiDiff) attention. (a) refers to the attention distraction in mainstream Transformer-based SNNs, where models excessively focus on irrelevant background information, leading to misclassification. (b) denotes the retinal lateral inhibition mechanism, illustrating how horizontal cells regulate neural responses to enhance contrast and suppress noise. (c) presents the architecture of SpiLiFormer. (d) and (e) represent the information processing flows of FF-LiDiff and FB-LiDiff attention, respectively.
@@ 3.4 Feedback-pathway Lateral Differential Inhibition Attention
The second forward propagation begins by taking the processed forward representation $S_{\text{stage3}}^{Ds}$ and the corresponding feedback information $X_{\text{FB}}^{1}$ as input to the first FB-LiDiff attention. The objective is to reduce model energy consumption by eliminating redundant computations in the non-feedback loop (i.e., Stage 1 and Stage 2). A detailed energy consumption analysis can be found in Appendix A.3. Additionally, during the second forward propagation, we share all parameter weights across the FB-LiDiff blocks to minimize the number of model parameters. Given the input $X_{2}^{m}$, the process described above can be formulated as:
@@ 3.5 Training Strategy and Loss Function
where $\alpha$ is a hyperparameter that dynamically balances the losses from two forward propagations. Unless explicitly specified otherwise, $\alpha$ is set to 0.5 in the subsequent experiments. The choice of this value is validated through an ablation study, as detailed in Appendix A.4. Additionally, $y$ represents the true label; $O_{\text{stage3}}^{N_{3}}$ and $(O_{\text{stage3}}^{N_{3}})^{{}^{\prime}}$ denote the outputs from the first and second forward propagations of the last FB-LiDiff block, respectively; $CH(\cdot)$ represents the classification head module, as described in previous studies [36, 35]; and $\mathcal{L}_{CE}$ denotes the cross-entropy loss function.
@@ 4 Experiments
Datasets Methods Architecture Param(M) Time Step Top-1 Acc(%)
@@ 4.5 Ablation Study
We conduct an ablation study on CIFAR-10 (static) and CIFAR10-DVS (neuromorphic) to assess the impact of FF-LiDiff and FB-LiDiff attention on model performance and inference latency. All ablation experiments follow the training details in Sec. 4.2 and Sec. 4.3, unless otherwise specified.
@@
As shown in Tab. 3, we observe that the model exhibits varying degrees of performance degradation on both datasets under the two aforementioned scenarios. Specifically, performance exhibits a decrease of 0.31% and 0.62% on the CIFAR-10 dataset, and 1.6% and 3.0% on the CIFAR10-DVS dataset. These results highlight the effectiveness of FB-LiDiff in mitigating attention distraction and improving model performance. However, its additional forward pass introduces higher inference latency, particularly on static datasets, with a 21–23% increase as shown in Tab. 5 in the appendix.
Datasets Methods Top-1 Acc(%)
@@ 5 Conclusion
6 Acknowledgments
This research was supported in part by Zhejiang Provincial Natural Science Foundation of China under the Grant No. LD25F020006.
@@ A.3 Energy Consumption Calculation of SNNs and ANNs
A.4 Selection of the Optimal $\alpha$ Hyperparameter
We perform an ablation study on both static and dynamic datasets to select the optimal value of the hyperparameter $\alpha$, as shown in Tab. 4. The results show that the model achieves peak accuracy when $\alpha=0.5$, while other values lead to varying degrees of performance degradation. As a result, we set $\alpha$ to 0.5 by default in all subsequent experiments.
Datasets $\alpha$ Value
0.3 0.4 0.5 0.6 0.7
CIFAR-10 96.35 96.40 96.63 96.36 96.16
CIFAR10-DVS 86.3 86.4 86.7 86.1 85.6
Table 4: Ablation Study of the $\alpha$ hyperparameter
A.5 Evaluation of Inference Latency
Datasets Inference Time per Sample (ms)
w/o FB-LiDiff w/ FB-LiDiff
CIFAR-10 0.7657 0.9322 (+21.7%)
CIFAR-100 0.6965 0.8581 (+23.2%)
CIFAR10-DVS 9.9433 10.0657 (+1.2%)
N-Caltech101 18.4273 19.8113 (+7.5%)
ImageNet-1K 58.9901 66.3027 (+12.4%)
Table 5: Inference time per sample (ms) across datasets.
We conduct a comprehensive evaluation of the inference latency introduced by FB-LiDiff due to its additional forward pass. As shown in Tab. 5, FB-LiDiff increases inference time by 1.2% to 23.2%, with over 20% overhead observed on static CIFAR datasets.
A.6 Supplementary Tables and Figures